OpenStudy (anonymous):

Write the equation of the line passes through (-1,4) with a slope of -1 in standard form

OpenStudy (anonymous):

standard form is y = mx + b where m is the slope y = 4 x = -1 m = -1 find b

OpenStudy (anonymous):

Standard form:Ax + By = C, y=mx+b is slope intercept form

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

so can you help?

OpenStudy (anonymous):

solve for the slope intercept form anyway, then change it to the standard form

OpenStudy (texaschic101):

y - y1 = m(x - x1) slope(m) = -1 x1 = -1 y1 = 4 y - 4 = -1(x - (-1) y - 4 = -x + 1 y = -x + 1 + 4 y = -x + 5 <-- slope intercept form Ax + By = C y = -x + 5 x + y = 5 < -- standard form